OpenStudy (anonymous):

3. As a valet at the mall, Davis works 7-hour shifts. He earns $4 per hour and keeps 70% of his tip money. How much does Davis need to be tipped so that he takes home exactly $87.50, before taxes are deducted, per shift? A $74.38 B $125.00 C $85.00 D $198.33

OpenStudy (ash2326):

7 hour shift so 7*4$=$28 he earns a day suppose he gets $x per day 70% of x=70/100 * x=0.7 x so 0.7x+28=87.5 0.7x=59.5 x= $85 he need to be tipped $ 85